Roulette: There 40 slots in a roulette wheel, 19 are red, 19 are black, and 2 are green. If you place a $1 bet on red and win, you get $2 (your original $1 and an additional $1) returned. What is the expected value of a $1.00 bet on red.
Answer #1

P(Red) = 19/40 = 0.475

So P(Not red) = 1 - 0.475 = 0.525

Therefore,

Expected value of $1.00 bet on red

= $1 * 0.475 - $1 * 0.525

= - $0.05